OpenStudy (anonymous):

Verify whether the given ordered pair is a solution to the system. If it is not a solution,explain why not? a.(-15.6,0.2) b.(-4,23) c.(2,12.3) y=47+3x y=15-2x y=4.5+5x y=8+0.5X y=12+x y=2.3+5x

OpenStudy (anonymous):

Verify whether the given ordered pair is a solution to the system. If it is not a solution,explain why not? a.(-15.6,0.2) b.(-4,23) c.(2,12.3) y=47+3x y=15-2x y=4.5+5x y=8+0.5X y=12+x y=2.3+5x

OpenStudy (istim):

Pardon me if I sound rude, but is it possible for you to just sub in those values and attempt to solve for....something??

OpenStudy (anonymous):

k

OpenStudy (istim):

Dunno if that does much...